OpenStudy (anonymous):

If (2+i)(2+2i)(2+3i)...............(2+ni)=x+iy then find the value of 5.8.13. ........(4+n^2).

ganeshie8 (ganeshie8):

\[|z_1z_2| = |z_1||z_2|\]

OpenStudy (anonymous):

you mean \[\large \bf \left| z_1+z_2 \right|=\left| z_1 \right|+\left| z_2 \right|\]

OpenStudy (anonymous):

right ??

OpenStudy (dan815):

Productsumthing n=1 to N of(C+nX)

OpenStudy (dan815):